Office Expense. The monthly phone expense for an office is $1094 less than the janitorial expense. Three times the janitorial expense minus four times the phone expense is $248. What is the total of the two expenses?


Answer by orca(409)

Let x and y be the phone expense and janitorial expense respectively.
As the phone expense for an office is $1094 less than the janitorial expense, we have:
y - x = 1094 ...........(1)
As three times the janitorial expense minus four times the phone expense is $248, we have
3y - 4x = 248 ..........(2)
From equation(1), we get y = x + 1094.
Substituting it into equation (2), we have
3(x + 1094) - 4x = 248
Solving for x, we obtain:
3x + 3282 - 4x = 248
-x = -3034
x = 3034
Substituting x = 3034 into y = x + 1094, we obtain
y = x + 1094 = 4138
So the total of the two expenses is x + y = 3034 + 4138 = $7172.